i have a square that is 34 inches what is the diagnal ??
thank you

Answer by unlockmath(1688)

Hello,
Use this formula:
C^2=A^2+B^2
C is the diagonal
C^2=34^2+34^2
C^2=2312
Square root both sides:
Diagonal is 48.08 inches (rounded off)
Make sense?
RJ
www.math-unlock.com